# Thermodynamic Consistent Description of Compact Stars of Two Interacting Fluids: The Case of Neutron Stars with Higgs Portal Dark Matter

### v1

@@ 3 The Higgs Portal Fermionic Dark Matter Model
$p_{\text{tot}}=$ $\ p_{\text{baryons}}+p_{\text{\rm DM}}+p_{\text{mesons+Higgs}}+p_{\text{leptons}}\,,$ (3.23)
$p_{\text{baryons}}=$ $\ \frac{\gamma}{3(2\pi)^{3}}\int_{0}^{k_{F}}\frac{k^{4}dk}{\sqrt{k^{2}+(M_{\rm N}^{*})^{2}}}\,,$ (3.24)
$p_{\text{\rm DM}}=$ $\ \frac{\gamma}{3(2\pi)^{3}}\int_{0}^{k_{F}^{\text{\rm DM}}}\frac{k^{4}dk}{\sqrt{k^{2}+(M_{\chi}^{*})^{2}}}\,,$ (3.25)
$p_{\text{mesons+Higgs}}=$ $\ -\left(\frac{1}{2}m_{s}^{2}\phi_{0}^{2}+\frac{1}{3}g_{2}\phi_{0}^{3}+\frac{1}{4}g_{3}\phi_{0}^{4}-\frac{1}{2}m_{V}^{2}V_{0}^{2}-\frac{1}{2}m_{\rho}^{2}b_{0}^{2}+\frac{1}{2}M_{h}^{2}h_{0}^{2}\right)\,,$ (3.26)
$p_{\text{leptons}}=$ $\ \sum_{l=e,\mu}\frac{1}{3\pi^{2}}\int_{0}^{k_{l}}\frac{k^{4}dk}{\sqrt{k^{2}+m_{l}^{2}}}\,.$ (3.27)
@@
$\mu_{n}$ $=\mu_{p}+\mu_{e}\,,\quad\mu_{e}=\mu_{\mu}\,,$ (3.28)
$\rho_{p}$ $=\rho_{e}+\rho_{\mu}\,,$ (3.29)
@@ 4 Results
Figure 2: Speed of sound versus nucleon and DM chemical potentials i.e. $\mu_{\rm N}$ and $\mu_{\chi}$ Each row is for a specific DM particle mass, that is, $M_{\chi}=100~{}$MeV (top left), $M_{\chi}=500~{}$MeV (top right), $M_{\chi}=1000~{}$MeV (bottom left) and $M_{\chi}=5000~{}$MeV (bottom right). We normalize the DM and nucleon chemical potentials with respect to the nucleon mass.
@@ 4.1 Radial Profiles and Mass-Radius Relation
By solving the structure equations for the chemical potentials, we can extract the radial profiles of the chemical potentials of the dark and visible sectors, as well as the DM and nucleon number densities ($N_{\rm DM}$ and $N_{\rm N}$, respectively) and the total mass of star. These are shown in the different panels of Figs. 3 and 4, where the results of the different columns correspond to DM particles masses of $M_{\chi}=100~{}$MeV, $M_{\chi}=500~{}$MeV, $M_{\chi}=1000$ MeV and $M_{\chi}=5000~{}$MeV. Note that the blue lines denote the DM and nucleonic matter mixed region and the red lines indicate the region where only nucleonic matter is present. These lines result from choosing two initial values for the chemical potentials for DM and nucleons, that is, $\mu_{\rm N}=M_{\rm N}+1.8$ and $\mu_{\rm DM}=M_{\chi}+0.1$ (in nucleon mass units).
As it can be seen in these figures, $\mu_{\chi}$ and $\mu_{\rm N}$ start from the mentioned fixed central values, vary with the distance to the center of the star and become constant when they reach $M_{\chi}$ and $M_{\rm N}$, respectively. In all cases shown, we obtain a DM core surrounded by nucleonic matter. The size of this core depends on the DM particle mass and the relation between the central chemical potentials. In particular, the larger the DM particle mass, the smaller the DM core. Note that if we had explored much larger values for the DM central chemical potential, we would have reached the situation where a DM halo is formed, as we will discuss later.
@@
With regard to the mass-radius relation, these are displayed in Fig. 5 for different DM particle masses and different central chemical potential ratios (denoted by the different coloured lines in each panel). The black curves in Fig. 5 for all DM particle masses indicate pure nucleonic matter. The other colored curves show admixture of DM with nucleons with different values of nonvanishing chemical potentials. For DM particle masses below the nucleon mass in the top panels of Fig. 5, we observe that the mass-radius relations deviate from the usual neutron star mass-radius configurations as a dark halo appears, exceeding the matter radius. This can be seen in the case of the largest DM chemical potential by comparing the solid green lines for the total mass versus the visible radius with respect to the green dashed lines for the total mass versus the DM radius. For DM particle masses around the nucleon mass a small amount of DM can be sufficient to substantially change the mass-radius relation, as seen in the lower-left panel of Fig. 5. In this case, the increase in the DM central chemical potential leads to more compact DM admixed neutron stars. Also, we note that in this case a dark halo still appears but for masses smaller than a solar mass, that are not represented here. For larger DM particle masses in the lower-right panel of Fig. 5, we do not observe a variation of the mass-radius curves with increasing DM chemical potential, as larger values are needed in order to see any deviation.
We now turn to the fraction of DM with respect to ordinary matter for the maximum mass configurations for radii around $10-15$ km of Fig. 5. These fractions can be computed by means of Eq. (3.12) and Eq. (3.13). The values are shown in Table 2. As it can be seen, increasing the central DM chemical potential (going from blue to green curves) enhances the fraction of the DM number density with respect to the nucleonic one for DM particle masses smaller than the nucleon one since there can be a halo formation in these cases instead of a core. Previous studies of neutron stars with DM have found similar results using other DM models when the fraction of DM with respect to ordinary matter is small and there is a small or no interaction between the two fluids [16, 10, 13, 27].
@@
Figure 5: Mass-radius relation for DM-admixed neutron stars. From the top left panel to the bottom right one, we consider for DM particle mass $M_{\chi}=100~{}$MeV, $M_{\chi}=500~{}$MeV, $M_{\chi}=1000$ MeV and $M_{\chi}=5000~{}$MeV. The different colored lines display the mass-radius relation for various DM chemical potentials at the center (in terms of $M_{\chi}$). Whereas the solid curves are the mass-radius relations for the total mass versus the visible radius, the green dashed curves are examples of the total mass versus the radius of DM.
@@ 4.2 Tidal Deformability
where $R$ is the radius of the star. The parameter $λ\uplambda$ is crucial in gravitational wave astrophysics, especially for binary neutron star mergers, as it influences the late inspiral phase of the gravitational waveform [114, 47, 42, 115, 116]. To compute the tidal deformability, we have to calculate $k_{2}$. In order to do so, the following differential equation in the dimensionless form has to be solved simultaneously with the structure equations [114, 47, 50]:
$r^{\prime}\frac{dy(r^{\prime})}{dr^{\prime}}+y(r^{\prime})^{2}+y(r^{\prime})e^{\lambda(r^{\prime})}\left\{1+4\pi r^{\prime 2}\left[p^{\prime}(r^{\prime})-\epsilon^{\prime}(r^{\prime})\right]\right\}+r^{\prime 2}Q(r^{\prime})$ $=$ $0,$ (4.3)
@@
$Q(r^{\prime})$ $=$ $4\pi e^{\lambda(r^{\prime})}\left[5\epsilon^{\prime}(r^{\prime})+9p^{\prime}(r^{\prime})+\frac{\epsilon^{\prime}(r^{\prime})+p^{\prime}(r^{\prime})}{dp^{\prime}/d\epsilon^{\prime}}\right]-6\frac{e^{\lambda(r^{\prime})}}{r^{\prime 2}}-\left[\frac{d\nu(r^{\prime})}{dr^{\prime}}\right]^{2}\,,$ (4.4)
@@
$e^{\lambda(r^{\prime})}$ $=$ $\left[1-\frac{2M^{\prime}(r^{\prime})}{r^{\prime}}\right]^{-1},$ (4.5)
$\frac{d\nu}{dr^{\prime}}$ $=$ $\frac{1}{r^{\prime}}\left[\frac{M^{\prime}(r^{\prime})+4\pi p^{\prime}(r^{\prime})r^{\prime 3}}{r^{\prime}-2M^{\prime}(r^{\prime})}\right].$
@@
$k_{2}$ $=$ $\frac{8C^{5}}{5}(1-2C)^{2}[2+2C(y-1)-y]$ (4.6)
$\times$ $\left\{2C[6-3y+3C(5y-8)]+4C^{3}[13-11y+C(3y-2)+2C^{2}(1+y)]\right.$
$\left.+3(1-2C)^{2}[2-y+2C(y-1)]\ln(1-2C)\right\}^{-1}\,,$
@@
$C=\frac{M(R)}{R}\,.$ (4.7)
The tidal deformability is then given by Eq. (4.2), whereas the dimesionless tidal deformability reads  [^4] The effective tidal deformability parameter $\tilde{\Lambda}$ for a binary neutron star merger with separate tidal deformability parameters $\Lambda_{1}$ and $\Lambda_{2}$ is given by [42] $\tilde{\Lambda}$ $=$ $\frac{16(m_{1}+12m_{2})m_{1}^{4}\Lambda_{1}+(m_{2}+12m_{1})m_{2}^{4}\Lambda_{2}}{13(m_{1}+m_{2})^{5}}.$ (4.8)
$\Lambda=\frac{2k_{2}}{3C^{5}}\,.$ (4.9)
@@
Compared to the limits extracted from GW170817 measurement of $70\lesssim\Lambda_{1.4}\lesssim 580$ [43], we find that our results for the dimensionless tidal deformability lie below the upper bound. The overall behaviour of tidal deformability in the presence of DM has been also studied in Refs. [7, 8, 12, 3, 16, 13, 31, 24, 11, 22, 28, 30, 29], but not finding the nonlinear behaviour of the tidal deformability with the total mass which we find for DM particle masses below the nucleon mass.
@@ 5 Summary and Conclusions
We have also estimated the tidal deformability for neutron stars with varying the DM content, which can be important for gravitational wave analyses, such as the event GW170817 observed by the LIGO and Virgo collaborations [119]. An increase in the central chemical potential of DM reduces the tidal deformability, making neutron stars less deformable under tidal forces as long as the dark halo has not formed. For regions where we have a large dark halo like for DM particle masses $100$ MeV and $500$ MeV the total tidal deformability of the star increases. These changes would influence the gravitational wave signal during the inspiral phase of mergers, as studied before [48, 114]. The alteration in the tidal deformability values has observational implications, as gravitational wave measurements can constrain the EoS of compact star matter such that an unusual value of the tidal deformability can signal the presence of DM which otherwise could not be reconciled with theoretical nuclear EoS models.
It is worth noting that our study has been focused on a specific DM model—the Higgs portal fermionic dark matter— in its mean-field approximation. While this provides valuable insights into the effects of DM on neutron star properties, further attempts are needed to explore other DM candidates and interactions. For instance, considering self-interacting DM, bosonic DM, or asymmetric DM models could reveal additional effects and lead to a more comprehensive understanding of the role of DM in neutron stars
@@ Acknowledgments
F.H. thanks Sina M.H. Hajkarim, Davoud Rafiei, Souroush Shakeri and Seokhoon Yun for useful discussions during this project. During the early stages of this project FH was partially supported by the research grant “New Theoretical Tools for Axion Cosmology” under the Supporting TAlent in ReSearch@University of Padova (STARS@UNIPD), Istituto Nazionale di Fisica nucleonice (INFN) through the Theoretical Astroparticle Physics (TAsP) project and the Deutsche Forschungsgemeinschaft (DFG) through the CRC-TR 211 project number 315477589-TRR 211. F.H. is supported by Homer Dodge postdoctoral fellowship and in part by DOE grant DE-SC0009956. F.H. also thank the organizers of workshop of Center for Theoretical Underground Physics and Related Areas (CETUP* - 2024), The Institute for Underground Science at Sanford Underground Research Facility (SURF), Lead, South Dakota for their hospitality and financial support. He also thanks the organizers of the Mitchell Conference in May 2024 at Texas A & M University for their hospitality and support during the initial stages of this project. L.T. acknowledges support from CEX2020-001058-M (Unidad de Excelencia “María de Maeztu") and PID2022-139427NB-I00 financed by the Spanish MCIN/AEI/10.13039/501100011033/FEDER,UE as well as from the Generalitat de Catalunya under contract 2021 SGR 171 and from the Generalitat Valenciana under contract CIPROM/2023/59. L.T. and J.S.B. acknowledge support by the CRC-TR 211 ’Strong-interaction matter under extreme conditions’- project Nr. 315477589 - TRR 211.

### v2

@@ 3 The Higgs Portal Fermionic Dark Matter Model
$\displaystyle p_{\text{tot}}=$ $\displaystyle\ p_{\text{baryons}}+p_{\text{\rm DM}}+p_{\text{mesons+Higgs}}+p_{\text{leptons}}\,$, (3.23)
$\displaystyle p_{\text{baryons}}=$ $\displaystyle\ \frac{\gamma}{3(2\pi)^{3}}\int_{0}^{k_{F}}\frac{k^{4}dk}{\sqrt{k^{2}+(M_{\rm N}^{*})^{2}}}\,$, (3.24)
$\displaystyle p_{\text{\rm DM}}=$ $\displaystyle\ \frac{\gamma}{3(2\pi)^{3}}\int_{0}^{k_{F}^{\text{\rm DM}}}\frac{k^{4}dk}{\sqrt{k^{2}+(M_{\chi}^{*})^{2}}}\,$, (3.25)
$\displaystyle p_{\text{mesons+Higgs}}=$ $\displaystyle\ -\left(\frac{1}{2}m_{s}^{2}\phi_{0}^{2}+\frac{1}{3}g_{2}\phi_{0}^{3}+\frac{1}{4}g_{3}\phi_{0}^{4}-\frac{1}{2}m_{V}^{2}V_{0}^{2}-\frac{1}{2}m_{\rho}^{2}b_{0}^{2}+\frac{1}{2}M_{h}^{2}h_{0}^{2}\right)\,$, (3.26)
$\displaystyle p_{\text{leptons}}=$ $\displaystyle\ \sum_{l=e,\mu}\frac{1}{3\pi^{2}}\int_{0}^{k_{l}}\frac{k^{4}dk}{\sqrt{k^{2}+m_{l}^{2}}}\,.$ (3.27)
@@
$\displaystyle\mu_{n}$ $\displaystyle=\mu_{p}+\mu_{e}\,,\quad\mu_{e}=\mu_{\mu}\,$, (3.28)
$\displaystyle\rho_{p}$ $\displaystyle=\rho_{e}+\rho_{\mu}\,$, (3.29)
@@ 4 Results
Figure 2: Speed of sound versus nucleon and DM chemical potentials i.e. $\mu_{\rm N}$ and $\mu_{\chi}$ . Each row is for a specific DM particle mass, that is, $M_{\chi}=100$~MeV (top left), $M_{\chi}=500$~MeV (top right), $M_{\chi}=1000$~MeV (bottom left) and $M_{\chi}=5000$~MeV (bottom right). We normalize the DM and nucleon chemical potentials with respect to the nucleon mass.
@@ 4.1 Radial Profiles and Mass-Radius Relation
By solving the structure equations for the chemical potentials, we can extract the radial profiles of the chemical potentials of the dark and visible sectors, as well as the DM and nucleon number densities ($N_{\rm DM}$ and $N_{\rm N}$, respectively) and the total mass of star. These are shown in the different panels of Figs. 3 and 4, where the results of the different columns correspond to DM particles masses of $M_{\chi}=100$~MeV, $M_{\chi}=500$~MeV, $M_{\chi}=1000$ MeV and $M_{\chi}=5000$~MeV. Note that the blue lines denote the DM and nucleonic matter mixed region and the red lines indicate the region where only nucleonic matter is present. These lines result from choosing two initial values for the chemical potentials for DM and nucleons, that is, $\mu_{\rm N}=M_{\rm N}+1.8$ and $\mu_{\rm DM}=M_{\chi}+0.1$ (in nucleon mass units)  [^4] To calculate the structure equations wherever we need the interpolation of EoS we use a machine learning technique called Random Forest algorithm to have a precise estimation of interpolated functions..
As it can be seen in these figures, $\mu_{\chi}$ and $\mu_{\rm N}$ start from the mentioned fixed central values, vary with the distance to the center of the star and become constant when they reach $M_{\chi}$ and $M_{\rm N}$, respectively. In all cases shown, we obtain a DM core surrounded by nucleonic matter. The size of this core depends on the DM particle mass and the relation between the central chemical potentials. In particular, the larger the DM particle mass, the smaller the DM core. Note that if we had explored much larger values for the DM central chemical potential, we would have reached the situation where a DM halo is formed, as we will discuss later. The central chemical potentials ($\mu_{\rm N}=M_{\rm N}+1.8$, $\mu_{\rm DM}=M_{\chi}+0.1$, in nucleon mass units) are chosen to explore a realistic range for neutron star cores ($\mu_{\rm N}\sim 1.5-2\,M_{\rm N}$) and a perturbative DM contribution, reflecting possible DM accumulation scenarios. These values ensure a DM core, while larger $\mu_{\rm DM}$ leads to halo formation (see Figs. 5 and 6). Masses and radii are presented in physical units (solar masses and km) in Figs. 5 and 6, converted from dimensionless solutions, with scaling independent of $M_{\chi}$ to facilitate comparison.
@@
With regard to the mass-radius relation, these are displayed in Fig. 5 for different DM particle masses and different central chemical potential ratios (denoted by the different coloured lines in each panel). The black curves in Fig. 5 for all DM particle masses indicate pure nucleonic matter. The other colored curves show admixture of DM with nucleons with different values of nonvanishing chemical potentials. For DM particle masses below the nucleon mass in the top panels of Fig. 5, we observe that the mass-radius relations deviate from the usual neutron star mass-radius configurations as a dark halo appears, exceeding the matter radius. This can be seen in the case of the largest DM chemical potential by comparing the solid green lines for the total mass versus the total radius with respect to the green dashed lines for the total mass versus the DM radius. For DM particle masses around the nucleon mass a small amount of DM can be sufficient to substantially change the mass-radius relation, as seen in the lower-left panel of Fig. 5. In this case, the increase in the DM central chemical potential leads to more compact DM admixed neutron stars. Also, we note that in this case a dark halo still appears but for masses smaller than a solar mass, that are not represented here. For larger DM particle masses in the lower-right panel of Fig. 5, we do not observe a variation of the mass-radius curves with increasing DM chemical potential, as larger values are needed in order to see any deviation.
We now turn to the fraction of DM with respect to ordinary matter for the maximum mass configurations for radii around $10-15$ km of Fig. 5. These fractions can be computed by means of Eq. (3.12) and Eq. (3.13). For neutron star matter (black curves), the maximum TOV mass is $\sim 2.7\,M_{\odot}$, consistent with observations. DM admixture softens the EoS, reducing the maximum mass to $1.8-2.0\,M_{\odot}$ for $M_{\chi}=1000\,\text{MeV}$ (lower-left panel), still within observational bounds. For lighter DM ($M_{\chi}=100\,\text{MeV}$), high $\mu_{\rm DM}$ (green curves) yields masses up to $2.5\,M_{\odot}$ due to halo formation, potentially exceeding current limits. If one finds a neutron star with mass above 3 or more solar masses it could be explained with the presence of DM. The values are shown in Table 2. As it can be seen, increasing the central DM chemical potential (going from blue to green curves) enhances the fraction of the DM number density with respect to the nucleonic one for DM particle masses smaller than the nucleon one since there can be a halo formation in these cases instead of a core. Previous studies of neutron stars with DM have found similar results using other DM models when the fraction of DM with respect to ordinary matter is small and there is a small or no interaction between the two fluids [16, 10, 13, 28].
Note that our current model focuses on the core EoS and omits DM effects in the crust due to limited understanding of Higgs portal DM interactions in this regime. While the crust contributes minimally to the total mass and radius for stars with mass above one solar mass, its inclusion could influence continuous GW signals from non-radial oscillations, as these depend on crust-core interactions. A detailed crust EoS with DM is needed to quantify continuous GW amplitudes, an avenue for future study.
@@
Figure 5: Mass-radius relation for DM-admixed neutron stars. From the top left panel to the bottom right one, we consider for DM particle mass $M_{\chi}=100$~MeV, $M_{\chi}=500$~MeV, $M_{\chi}=1000$ MeV and $M_{\chi}=5000$~MeV. The different colored lines display the mass-radius relation for various DM chemical potentials at the center (in terms of $M_{\chi}$). Whereas the solid curves are the mass-radius relations for the total mass versus the total radius, the green dashed curves are examples of the total mass versus the radius of DM.
@@ 4.2 Tidal Deformability
[…]
$\displaystyle r^{\prime}\frac{dy(r^{\prime})}{dr^{\prime}}+y(r^{\prime})^{2}+y(r^{\prime})e^{\lambda(r^{\prime})}\left\{1+4\pi r^{\prime 2}\left[p^{\prime}(r^{\prime})-\epsilon^{\prime}(r^{\prime})\right]\right\}+r^{\prime 2}Q(r^{\prime})$ $\displaystyle=$ $\displaystyle 0,$ (4.3)
@@
$\displaystyle Q(r^{\prime})$ $\displaystyle=$ $\displaystyle 4\pi e^{\lambda(r^{\prime})}\left[5\epsilon^{\prime}(r^{\prime})+9p^{\prime}(r^{\prime})+\frac{\epsilon^{\prime}(r^{\prime})+p^{\prime}(r^{\prime})}{dp^{\prime}/d\epsilon^{\prime}}\right]-6\frac{e^{\lambda(r^{\prime})}}{r^{\prime 2}}-\left[\frac{d\nu(r^{\prime})}{dr^{\prime}}\right]^{2}\,$, (4.4)
@@
$\displaystyle e^{\lambda(r^{\prime})}$ $\displaystyle=$ $\displaystyle\left[1-\frac{2M^{\prime}(r^{\prime})}{r^{\prime}}\right]^{-1},$ (4.5)
$\displaystyle\frac{d\nu}{dr^{\prime}}$ $\displaystyle=$ $\displaystyle\frac{1}{r^{\prime}}\left[\frac{M^{\prime}(r^{\prime})+4\pi p^{\prime}(r^{\prime})r^{\prime 3}}{r^{\prime}-2M^{\prime}(r^{\prime})}\right].$
@@
$\displaystyle k_{2}$ $\displaystyle=$ $\displaystyle\frac{8C^{5}}{5}(1-2C)^{2}[2+2C(y-1)-y]$ (4.6)
$\displaystyle\times$ $\displaystyle\left\{2C[6-3y+3C(5y-8)]+4C^{3}[13-11y+C(3y-2)+2C^{2}(1+y)]\right$.
$\displaystyle\left.+3(1-2C)^{2}[2-y+2C(y-1)]\ln(1-2C)\right\}^{-1}\,$,
@@
$\displaystyle C=\frac{M(R)}{R}\,.$ (4.7)
The tidal deformability is then given by Eq. (4.2), whereas the dimensionless tidal deformability reads  [^5] The effective tidal deformability parameter $\tilde{\Lambda}$ for a binary neutron star merger with separate tidal deformability parameters $\Lambda_{1}$ and $\Lambda_{2}$ is given by [47] $\displaystyle\tilde{\Lambda}$ $\displaystyle=$ $\displaystyle\frac{16(m_{1}+12m_{2})m_{1}^{4}\Lambda_{1}+(m_{2}+12m_{1})m_{2}^{4}\Lambda_{2}}{13(m_{1}+m_{2})^{5}}.$ (4.8)
$\displaystyle\Lambda=\frac{2k_{2}}{3C^{5}}\,.$ (4.9)
@@
Compared to the limits extracted from GW170817 measurement of $70\lesssim\Lambda_{1.4}\lesssim 580$ [48], we find that our results for the dimensionless tidal deformability lie below the upper bound. The overall behaviour of tidal deformability in the presence of DM has been also studied in Refs. [7, 12, 3, 16, 13, 32, 25, 11, 23, 29, 31, 30], but not finding the nonlinear behaviour of the tidal deformability with the total mass which we find for DM particle masses below the nucleon mass.
@@ 5 Summary and Conclusions
We have also estimated the tidal deformability for neutron stars with varying the DM content, which can be important for gravitational wave analyses, such as the event GW170817 observed by the LIGO and Virgo collaborations [122]. An increase in the central chemical potential of DM reduces the tidal deformability, making neutron stars less deformable under tidal forces as long as the dark halo has not formed. For regions where we have a large dark halo like for DM particle masses $100$ MeV and $500$ MeV the total tidal deformability of the star increases. These changes would influence the gravitational wave signal during the inspiral phase of mergers, as studied before [53, 117]. The alteration in the tidal deformability values has observational implications, as gravitational wave measurements can constrain the EoS of compact star matter such that an unusual value of the tidal deformability can signal the presence of DM which otherwise could not be reconciled with theoretical nuclear EoS models. Observational constraints on DM halos around neutron stars remain limited. LIGO (GW170817) and NICER data constrain total mass and radius but do not directly probe DM distributions. Indirectly, pulsar timing or continuous GW signals could reveal DM halo effects via altered rotational dynamics or oscillation modes [49, 50], though no definitive evidence exists yet. For $M_{\chi}=100\,\text{MeV}$ with large $\mu_{\rm DM}$, our predicted halos (Fig. 5) extend radii beyond typical neutron star sizes ($R>15\,\text{km}$), potentially detectable with future high-precision radius measurements (e.g., from X-ray timing) [17].
It is worth noting that our study has been focused on a specific DM model–the Higgs portal fermionic dark matter– in its mean-field approximation. While this provides valuable insights into the effects of DM on neutron star properties, further attempts are needed to explore other DM candidates and interactions. For instance, considering self-interacting DM, bosonic DM, or asymmetric DM models could reveal additional effects and lead to a more comprehensive understanding of the role of DM in neutron stars.
@@ Acknowledgments
F.H. thanks Raghuveer Garani, Sina M.H. Hajkarim, Davoud Rafiei, Souroush Shakeri and Seokhoon Yun for useful discussions during this project. During the early stages of this project FH was partially supported by the research grant “New Theoretical Tools for Axion Cosmology” under the Supporting TAlent in ReSearch@University of Padova (STARS@UNIPD), Istituto Nazionale di Fisica nucleonice (INFN) through the Theoretical Astroparticle Physics (TAsP) project and the Deutsche Forschungsgemeinschaft (DFG) through the CRC-TR 211 project number 315477589-TRR 211. F.H. is supported by Homer Dodge postdoctoral fellowship and in part by DOE grant DE-SC0009956. F.H. also thanks the organizers of workshop of Center for Theoretical Underground Physics and Related Areas (CETUP* - 2024), The Institute for Underground Science at Sanford Underground Research Facility (SURF), Lead, South Dakota for their hospitality and financial support. Moreover, he thanks the organizers of the Mitchell Conference in May 2024 at Texas A & M University for their hospitality and support during the initial stages of this project. L.T. acknowledges support from CEX2020-001058-M (Unidad de Excelencia “María de Maeztu") and PID2022-139427NB-I00 financed by the Spanish MCIN/AEI/10.13039/501100011033/FEDER,UE as well as from the Generalitat de Catalunya under contract 2021 SGR 171 and from the Generalitat Valenciana under contract CIPROM/2023/59. L.T. and J.S.B. acknowledge support by the CRC-TR 211 ’Strong-interaction matter under extreme conditions’- project Nr. 315477589 - TRR 211.